OpenStudy (anonymous):

Find the exact value of the following: a) sin^2 (75^0) / sin45^0 cos^2 (30^0) b) tan(135^0) + sin30^0 x cot120^0

OpenStudy (anonymous):

\[a=[\sin(30+45)]/[\sin45\times \cos^230]\] \[=[\sin30 \cos45 +\cos30 \sin45]^2/[\sin45\times \cos^230]\] then just plug in the numbers b tan(135)=tan(90+45)\[=\sin(90+45)/\cos(90+45)\] cot(120)=cot(90+30) try to use this hint to find the answer